If the coefficient of friction between tires and roadway is 0.50, what is the smallest radius at which


a car can turn on a horizontal road when its speed is 30 miles/hour?

1
Expert's answer
2022-05-31T10:04:58-0400
"ma=F""\frac{mv^2}{R}=\mu mg"

"R=\frac{v^2}{g}=\frac{(13.4\:\rm m/s)^2}{9.8\:\rm m/s^2}=18.3\:\rm m"